How many Ne atoms are there in a 40.8 gram sample of Ne ?

How many grams of Ba are there in a sample of Ba that contains 8.26×1023atoms?

A sample of Hg weighs 31.4 grams. Will a sample of Fe that contains the same number of atoms weigh more or less than 31.4 grams? (more, less):

Calculate the mass of a sample of Fe that contains the same number of atoms.
Answer #1

a.

20g of Ne = 1 mole

20g of Ne   = 6.023*10^23 atoms

40.8g of Ne = 6.023*10^23*40.8/20   = 1.23*10^24 Ne atoms

b.

6.023*10^23 atoms = 1 mole of Ba

6.023*10^23 atoms = 137.327g of Ba

8.26*10^23 atoms   = 137.327*8.26*10^23/(6.023*10^23)   = 188.33g of Ba

c.

1 mole of Hg = 6.023*10^23 atoms

200.59g of Hg   = 6.023*10^23 atoms

31.4g of Hg = 6.023*10^23*31.4/200.59   = 9.43*10^22 atoms

6.23*10^23 atom = 55.875g of Fe

9.43*10^23 atom   = 55.875*9.43*10^23/(6.23*10^23) = 84.57g of Fe

more >>>answer